U . FUJIUgu.93 4. Solve for z in each of the following: a) 422–3 = 7.0 b) 4.82 = 7 · 32-1
Answer #1

a) 422-3 = 1 (2-3) log 42 = tog 7 - 2-3 2-3= log 2 Tagua. 2-3 = 0.5206 12=3.5206 I b) 4.8= 7132-1 => log (4.82) tag (7.32-4)